OpenStudy (anonymous):

add (-1/3)+(-2/7)+(4/9)

OpenStudy (anonymous):

@LoveYou*69 @Easyaspi314 @FutureMathProfessor

OpenStudy (anonymous):

MEDAL WILL BE GIVEN

OpenStudy (anonymous):

You have denominators of 3,7,9. You need the LCD. What number will be the LCD?

OpenStudy (anonymous):

63

OpenStudy (anonymous):

@mathnoobie10 @mathstudent55 @mathkid1

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Let's change all fractions to a denominator of 63.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

21/63 9/63 7/63

OpenStudy (anonymous):

1/3 = 21/63 2/7 = 18/63 4/9 = 28/63 So, you have -21/63 -18/63 + 28/63 = -11/63 Agree?

OpenStudy (anonymous):

how did you get 18/63 and 28/63 ????

OpenStudy (anonymous):

For 4/....9 into 63 goes 7 times, and 7 times 4 is 28.

OpenStudy (anonymous):

THANK YOU SO MUCH!!! GOD BLESS YOU!!

OpenStudy (anonymous):

welcome.